# Explicit Constructions of Sum-Rank Metric Codes from Quadratic Kummer Extensions

### v1

@@ Abstract
This paper presents new constructions of sum-rank metric codes derived from algebraic function fields, while the existing results about such codes remain limited. A central challenge in this field involves the determination of parameters. We address this challenge through quadratic Kummer extensions, and propose two general constructions of $2\times 2$ sum-rank codes. The novel sum-rank metric code constructions derived from algebraic function fields demonstrate superior code lengths compared to conventional coding paradigms, particularly when contrasted with linearized Reed-Solomon codes under equivalent code parameter constraints. We also establish explicit parameters including dimensions and minimum distances of our codes. Finally, an illustrative example through elliptic function fields is constructed for validating the theoretical framework.
@@ I Introduction
Algebraic function fields over finite fields provide a foundational framework for Hamming-metric code construction [15]. Unified methodological frameworks are constructed through exploitation of algebraic structures inherent in function fields, enabling simultaneous characterization of both minimum distance parameters and dual code invariants [41]. Specifically, a linear $[n,k]$ Hamming metric code $C$ over $\mathbb{F}_{q}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance
@@
where $wt_{H}$ counts non-zero coordinate positions. Classical algebraic geometry (AG) codes are constructed by evaluating functions in Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
For a linear $[n,k,d_{H}]$ code $C$, the Singleton bound asserts that
@@
When the equality holds, this code is called a maximum distance separable (MDS) code. Reed-Solomon (RS) codes [18], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent Hamming metric MDS codes. Their maximal length of $q+1$ over $\mathbb{F}_{q}$ coincides with the main conjecture of MDS codes.
Very recently, sum-rank metric codes have attracted significant attention due to applications in multishot network coding [22, 24, 31, 35], space-time coding [40], and distributed storage [23, 25]. For more applications and details, we refer the reader to [27]. The basic definitions are given as follows.
@@
Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, denote
$\Pi:=\Pi_{q}(m_{1}\times n_{1}|\cdots|m_{s}\times n_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{m_{i}\times n_{i}}$
where $\mathbb{F}_{q}^{m_{i}\times n_{i}}$ denotes the space of $m_{i}\times n_{i}$ matrices over $\mathbb{F}_{q}$. Note that $\Pi$ forms an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}m_{i}n_{i}$. The sum-rank weight of
@@
$wt_{sr}(X):=\sum\limits_{i=1}^{s}rank(X_{i}).$
The sum-rank distance between $X,Y\in\Pi$ is
@@
The sum-rank weight and distance induce a metric on $\Pi$. We formally define a linear sum-rank metric code as follows.
@@
A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$ with minimum sum-rank distance
$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in\Pi\}.$
For notational conciseness, we adopt the convention of referring to sum-rank metric codes as sum-rank codes throughout this work. The sum-rank code framework constitutes a natural generalization of classical coding systems: (i) When configured with block uniformity parameters $n_{i}=m_{i}=1$ for all $1\leq i\leq s$, these codes precisely recover the Hamming metric codes of length $s$ through coordinate-wise symbol comparison; (ii) In the monolithic regime where $s=1$, the construction collapses to the original rank-metric codes introduced by Delsarte [12]. This multifunctional parametrization reveals the intrinsic capacity of the codes to interpolate between symbol-level and matrix-level error protection strategies.
The theoretical landscape of sum-rank codes is underpinned by three foundational pillars as established in prior research:
  1. 1.
    Structural Properties: Code duality mechanisms and weight enumerator analysis [1, 5]
  2. 2.
    Distance Bounds: Fundamental limitations including Singleton-type and Johnson-type bounds [37]
  3. 3.
    Decoding Thresholds: Characterization of list-decoding capacity regions [2]
The Singleton bound for sum-rank codes was provided in [5] and [25]. For the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $10\leq\delta\leq n_{j}-$ 1. Then
@@
A code $C\subset\Pi$ attaining this bound is called maximum sum-rank distance (MSRD). For the special case $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound is of the form
@@
Throughout this paper, we focus on sum-rank metric codes with these uniform parameters.
Constructions of “good” sum-rank metric codes can be found in [1, 6, 7, 9, 23, 26, 28, 29, 30, 33, 34]. A standard technique constructs $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block. This ensures that the dimension $\kappa$ of $C$ over $\mathbb{F}_{q^{m}}$ satisfies $d_{sr}+\kappa\leq sn+1$, matching the Singleton bound for the Hamming metric. Moreover, efficient algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 16, 39].
One of the most prominent rank metric codes is the Gabidulin code, constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 14, 19, 21, 32]. Among sum-rank codes, linearized Reed-Solomon codes [23] represent the most extensively investigated families to date, see [8, 17, 20, 33, 38]. Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as rank metric and sum-rank metric analogues of RS codes.
@@
While existing algebraic geometric constructions [9, 28] for sum-rank metric codes are non-trivial, most results are constructed using towers of function fields as illustrated in Figure 1. In [4], E. Berardini and X. Caruso established the first geometric construction of sum-rank metric codes. Their approach employs the Ore polynomial ring [36] over an extension field derived from the completion of $F$ at specified places and subsequently develops the theory of Riemann-Roch spaces over this ring. This framework enables them to introduce linearized algebraic geometry codes and presents explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
A principal motivation for constructing codes via algebraic function fields lies in their ability to provide numerous evaluation places bounded by the Hasse-Weil bound with higher genus. Consequently, codes with long lengths can be constructed over small finite fields. Classical algebraic geometry codes are known to asymptotically surpass the Gilbert-Varshamov (GV) bound [43]. Recent work [37] demonstrates that random linear sum-rank metric codes attain the asymptotic GV-like bound with high probability. This raises a natural question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
@@
In contrast to the construction in [4], we propose an alternative method to construct sum-rank metric codes via function fields. The key tool of our construction is the local integral basis of a place in function field extensions. Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and $F^{\prime}/F$ a finite separable function field extension. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ forms a vector space over $\mathcal{O}_{P}$. Consequently, a function in $F^{\prime}/\mathbb{F}^{\prime}_{q}$ that is regular at the places lying over $P$ corresponds to a matrix over $F$, where $\mathbb{F}^{\prime}_{q}$ is a finite extension of $\mathbb{F}_{q}$. When $F^{\prime}/F$ is Galois, we establish a linearized-polynomial-like map on $\mathcal{O}^{\prime}_{P}$ through the direct sum of Riemann-Roch spaces, yielding sum-rank metric codes via evaluation of the function matrix at rational places. While our construction overlaps with [4], it differs substantially by emphasizing the intrinsic structure of function field extensions. By sacrificing certain subspaces in the direct sum of Riemann-Roch spaces, we explicitly determine the parameters of our sum-rank codes.
@@
The remainder of this paper is organized as follows. Section II introduces the foundational concepts necessary for our purpose. In Section III, we establish a general construction of sum-rank metric codes through function field extensions. Section IV presents the explicit sum-rank metric codes via quadratic Kummer extensions and introduces a new specialized construction, including a case of elliptic function fields. Section V concludes this paper.
@@ II Preliminaries and basic concepts
This section reviews some basic definitions and notations, primarily following [41].
@@ II-A Places and divisors
Let $q$ be a prime power. An algebraic function field $F/\mathbb{F}_{q}$ denotes a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$. A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$ with
$\deg(P)=[\mathcal{O}_{P}/P:\mathbb{F}_{q}].$
If $\deg(P)=1$, then $P$ is called rational. The prime element of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{p}$. Any non-zero $z\in F$ has a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$. The associated discrete valuation
@@
satisfies $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality of $v_{P}$ is characterized as follows:
@@
For any discrete valuation $v_{P}$,
@@
A divisor $G$ is defined as the formal sum of places
@@
with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}$. For any function $z\in F$,
@@
is called the principal divisor. Given two divisors $G_{1},G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places. The associated Riemann-Roch space is defined as
@@
The space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
@@
[41, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
@@ II-B Extensions of function fields
A finite algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $[F^{\prime}:F]<\infty$ and $m\geq 1$. A place $P^{\prime}$ of $F^{\prime}$ is said lying over $P$ if $P\subseteq P^{\prime}$, denoted $P^{\prime}|P$. The ramification index $e(P^{\prime}|P)$ is defined as the integer satisfying
@@
for all $x\in F$, while the relative degree of $P^{\prime}$ over $P$ is given by
@@
For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ with $P^{\prime}|P$. The conorm of $P$ is defined as
@@
Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ has the form
@@
[41, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$. Then
@@
If $[F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$. In this case, it follows that $e_{i}=f_{i}=1$. When $e(P^{\prime}|P)>1$ for some $P^{\prime}$, the place $P$ is ramified in $F^{\prime}/F$. Moreover, if there exists a place $P^{\prime}$ with $e(P^{\prime}|P)=n$, then $P$ is totally ramified in $F^{\prime}/F$.
II-C Galois Extensions and Kummer Extensions
An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension. Let $\text{Gal}(F^{\prime}/F)$ be the Galois group of $F^{\prime}/F$. For any place $P$ of $F$, the Galois group acts transitively on the set of places lying over $P$.
Kummer extensions constitute a special class of Galois extensions. We present their definition and key properties.
Definition II.1.
Let $\mathbb{F}_{q}$ contain a primitive $n$-th root of unity $\zeta$. Suppose that $u\in F$ satisfies
$u\neq\omega^{d}$
for all $\omega\in F$ and $1<d|n$. Let $F^{\prime}=F(y)$ with $y^{n}=u$ be an extension of $F/\mathbb{F}_{q}$. Then $F^{\prime}$ is called a Kummer extension of $F$ with:
  • $\text{Gal}(F^{\prime}/F)=\langle\sigma\rangle$ where $\sigma(y)=\zeta y$;
  • For any place $P$ of $F$ and $P^{\prime}$ lying over $P$, $e(P^{\prime}|P)=\frac{n}{\gcd(n,v_{P}(u))}$;
  • If there exists a place $Q$ of $F$ with $\gcd(n,v_{Q}(u))=1$, then $\mathbb{F}_{q}$ remains the constant field of $F^{\prime}$, and
    $g(F^{\prime})=1+n(g+1)+\frac{1}{2}\sum\limits_{P}(n-\gcd(n,v_{P}(u)))\deg(P);$
  • For any place $P$ with $v_{P}(u))\geq 0$, the basis $\{1,y,\ldots,y^{n-1}\}$ is a local integral basis of $F^{\prime}/F$ for $P$.

### v2

# Constructing Sum-Rank Metric Codes from Quadratic Galois Extensions of Function Fields

@@ Abstract
This paper introduces new constructions of sum-rank metric codes derived from algebraic function fields, as existing results on such codes remain limited. A major challenge lies in the determination of their parameters. We address this issue by employing quadratic Galois extensions, proposing two general constructions of $2\times 2$ sum-rank codes. Analogous to algebraic geometry codes in the Hamming metric, our codes achieve a larger block length compared to existing constructions. We determine explicit parameters including dimensions and minimum distances of our codes, and we present an illustrative example using elliptic function fields. Finally, we discuss the asymptotic behavior of our codes and compare them with the Gilbert-Varshamov-like bound for sum-rank metric codes.
@@ I Introduction
Algebraic function fields over finite fields provide a fundamental framework for modern error-correcting codes in the Hamming metric [16], as the function field perspective enables a deeper analysis of code parameters, including minimum distance and duality properties [43]. Specifically, a linear $[n,k]$ Hamming metric code $C$ over $\mathbb{F}_{q}$ is a $k$-dimensional subspace of $\mathbb{F}_{q}^{n}$ with minimum Hamming distance
@@
where $wt_{H}$ denotes the Hamming weight, counting non-zero coordinate positions. Classical algebraic geometry (AG) codes are constructed by evaluating functions from Riemann-Roch spaces at rational places, enabling explicit determination of code parameters.
For a linear $n,k,d_{H}]$ code $C$, the Singleton bound states that
@@
When equality holds, the code is called a maximum distance separable (MDS) code. Reed-Solomon (RS) codes [19], derived from rational function fields $\mathbb{F}_{q}(x)/\mathbb{F}_{q}$ with $x$ transcendental over $\mathbb{F}_{q}$, represent prominent examples of Hamming metric MDS codes. Their maximum length of $q+1$ over $\mathbb{F}_{q}$ aligns with the main conjecture of MDS codes.
Recently, sum-rank metric codes have garnered significant attention due to their applications in multishot network coding [24, 26, 33, 37], space-time coding [42], and distributed storage [25, 27]. For further details and applications, we refer the reader to [29]. The basic definitions are provided as follows.
@@
Given $s\geq 1$ and positive integers $n_{1},\ldots,n_{s},m_{1},\ldots,m_{s}$ with $n_{i}\leq m_{i}$ and $m_{1}\geq m_{2}\cdots\geq m_{s}$, define
$\Pi:=\Pi_{q}(n_{1}\times m_{1}|\cdots|n_{s}\times m_{s})=\bigoplus_{i=1}^{s}\mathbb{F}_{q}^{n_{i}\times m_{i}}$
where $\mathbb{F}_{q}^{n_{i}\times m_{i}}$ denotes the space of $n_{i}\times m_{i}$ matrices over $\mathbb{F}_{q}$. Note that $\Pi$ is an $\mathbb{F}_{q}$-linear space of dimension $\sum_{i=1}^{s}n_{i}m_{i}$. The sum-rank weight of
@@
$wt_{sr}(X):=\sum\limits_{i=1}^{s}{\rm rank}(X_{i}).$
The sum-rank distance between $X$ and $Y$ in $\Pi$ is
@@
The sum-rank weight and distance induce a metric on $\Pi$. We present the definition as follows.
@@
A linear sum-rank metric code $C_{sr}$ is an $\mathbb{F}_{q}$-linear subspace of $\Pi$. The minimum sum-rank distance of $C_{sr}$ is defined as
$d_{sr}(C):=\min_{X\neq Y}\{wt_{sr}(X-Y)|X,Y\in C_{sr}\}.$
Sum-rank metric codes generalize Hamming metric codes of length $s$ when $n_{i}=m_{i}=1$ for all $1\leq i\leq s$. When $s=1$, they reduce to rank metric codes, which were first introduced by Delsarte [12].
Fundamental properties and bounds for sum-rank metric codes are examined in [1, 2, 5, 39]. The Singleton bound was established in [5] and [27]. In the general case, let $j$ and $\delta$ be the unique integers satisfying $d-1=\sum_{i=1}^{j-1}n_{i}+\delta$ with $0\leq\delta\leq n_{j}-1$. Then
@@
A code $C\subseteq\Pi$ that attains this bound is called a maximum sum-rank distance (MSRD) code. For the special case where $n_{1}=\cdots=n_{s}=n$ and $m_{1}=\cdots=m_{s}=m$, this bound takes the form
@@
Throughout this paper, we concentrate on sum-rank metric codes with these uniform parameters.
Let $N:=sn$. For fixed positive real numbers $\mathcal{R}_{sr},\delta_{sr}\in(0,1)$ satisfying
$\mathcal{R}_{sr}<$ $\delta_{sr}^{2}-\delta_{sr}(2+\frac{2}{N})+1+\frac{2}{N}+\frac{1}{N^{2}}$
$-\frac{\sum_{i=1}^{d_{sr}N-1}\log_{q}(1+\frac{s-1}{i})+\log_{q}(\delta_{sr}N-1)}{Nm}-\frac{\log_{q}(\gamma_{q})}{nm},$
where $\gamma_{q}=\prod_{i=1}^{\infty}(1-q^{-i})^{-1}$, there exists a sum-rank code with rate at least $\mathcal{R}_{sr}$ and relative minimum sum-rank distance at least $\delta_{sr}$. When $m=\xi n$ goes to the infinity and $m\in\omega(\log_{q}(s))$, where $\xi$ is a constant, we have
$\mathcal{R}_{sr}\sim\delta_{sr}^{2}-\delta_{sr}(1+\frac{1}{\xi})+1.$
Recent work [39] demonstrates that random linear sum-rank metric codes attain the asymptotic Gilbert-Varshamov-like (GV-like) bound with high probability. Consider the case where $n=m$ and the number of blocks $s$ goes to infinity. Then the previous bound is asymptotically provided by [4]:
$\mathcal{R}_{sr}<$ $(\delta_{sr}-1)^{2}-\frac{\delta_{sr}}{n}\log_{q}(1+\frac{1}{\delta_{sr}n})-\frac{\log_{q}(1+\delta_{sr}n)}{n^{2}}$
$-\frac{\log_{q}(\gamma_{q})}{n^{2}}+o(1).$
Constructions of “good” sum-rank metric codes are available in several works [1, 6, 7, 9, 25, 28, 30, 31, 32, 35, 36]. A standard technique involves constructing $C$ as a classical Hamming metric code of length $sn$ over $\mathbb{F}_{q^{m}}$ and then applies rank metric code constructions to each $n$-length block. This ensures that the dimension $\kappa$ of $C$ over $\mathbb{F}_{q^{m}}$ satisfies $d_{sr}+\kappa\leq sn+1$, matching the Singleton bound for the Hamming metric. Moreover, algorithms over $\mathbb{F}_{q^{m}}$ can be employed for decoding sum-rank metric codes, see [3, 10, 17, 41].
One of the most prominent rank metric codes is the Gabidulin code, which is constructed using linearized polynomials over $\mathbb{F}_{q^{m}}$, see [13, 15, 21, 23, 34]. Among sum-rank metric codes, linearized Reed-Solomon codes [25] represent the most extensively investigated family to date, see [8, 18, 22, 35, 40]. Gabidulin codes and linearized RS codes are maximum rank distance (MRD) and MSRD respectively, serving as generalizations of RS codes.
@@
While non-trivial algebraic geometric constructions for sum-rank metric codes exist, see [9, 30], most results are constructed using towers of function fields as illustrated below. In [4], E. Berardini and X. Caruso established the first geometric construction of sum-rank metric codes. Their approach employs the Ore polynomial ring [38] over an extension field derived from the completion of $F$ at specified places, and subsequently develops the theory of Riemann-Roch spaces over this ring. This framework allows them to define linearized algebraic geometry codes and present explicit constructions of sum-rank metric codes with parameters consistent with classical algebraic geometry codes.
A principal motivation for constructing codes via algebraic function fields lies in the availability of numerous evaluation places bounded by the Hasse-Weil bound for higher genus. Consequently, codes with long lengths can be constructed over small finite fields. Classical algebraic geometry codes are known to asymptotically surpass the Gilbert-Varshamov (GV) bound [45]. This naturally raises the question: whether sum-rank codes derived from algebraic geometry also possess parameters surpassing the GV bound.
@@
In contrast to the construction in [4], we propose an approach for constructing sum-rank metric codes using function fields. The central element of our construction is the local integral basis at a place in function field extensions. Specifically, let $\mathcal{O}_{P}$ be the valuation ring at a place $P$ of $F/\mathbb{F}_{q}$, and let $F^{\prime}/F$ be a finite separable function field extension. The integral closure $\mathcal{O}^{\prime}_{P}$ of $\mathcal{O}_{P}$ in $F^{\prime}$ constitutes a vector space over $\mathcal{O}_{P}$. Consequently, a function in $F^{\prime}/\mathbb{F}^{\prime}_{q}$ that is regular at the places lying over $P$ corresponds to a matrix over $F$, where $\mathbb{F}^{\prime}_{q}$ is a finite extension of $\mathbb{F}_{q}$. When $F^{\prime}/F$ is Galois, we define a map on $\mathcal{O}^{\prime}_{P}$, via the direct sum of Riemann-Roch spaces. This yields sum-rank metric codes through the evaluation of the function matrix at rational places. Although our construction shares some similarities with [4], it significantly differs by focusing on the intrinsic structure of function field extensions. By decomposing the Riemann-Roch spaces as a direct sum of subspaces, we explicitly determine the parameters of our sum-rank codes. Furthermore, when $q$ is a special square, we analyze the asymptotic behavior of our codes and compare them with the GV-like bound.
@@
The remainder of this paper is organized as follows. Section II introduces the fundamental concepts required for our purpose. In Section III, we develop a general construction of sum-rank metric codes via function field extensions. Section IV provides the explicit sum-rank metric codes using quadratic Galois extensions and presents a new specialized construction, including an instance involving elliptic function fields and the comparison with the GV-like bound. Section V concludes this paper.
@@ II Preliminaries and basic concepts
This section reviews fundamental definitions and notations, primarily based on [43].
@@ II-A Places and divisors
Let $q$ be a prime power. An algebraic function field $F/\mathbb{F}_{q}$ is a finite extension of $\mathbb{F}_{q}(x)$ with genus $g(F)$, where $x$ is transcendental over $\mathbb{F}_{q}$. A place $P$ is the maximal ideal of a valuation ring $\mathcal{O}_{P}$ in $F/\mathbb{F}_{q}$. Denote by $\hat{F}_{P}:=\mathcal{O}_{P}/P$ the residue field of $P$, which is a finite extension of $\mathbb{F}_{q}$. The degree of $P$ is defined as
$\deg(P)=[\hat{F}_{P}:\mathbb{F}_{q}].$
If $\deg(P)=1$, then $P$ is called rational. The prime element (or uniformizing parameter) of $P$ is a function $t\in P$ such that $P=t\mathcal{O}_{P}$. Any non-zero element $z\in F$ admits a unique representation $z=t^{n}u$ with $u\in\mathcal{O}^{*}_{P}$. The discrete valuation associated with $P$ is
@@
, satisfying $v_{P}(z)=n$ and $v_{P}(0)=\infty$. The strict triangle inequality is characterized as follows:
@@
For any discrete valuation $v_{P}$ with associated place $P$,
@@
A divisor $G$ is defined as a formal sum of places
@@
with $n_{P}\in\mathbb{Z}$ and $\deg(G)=\sum\limits_{P}n_{P}\deg(P)$. The support of $G$ is defined as
$\text{supp}(G):=\{P\mid n_{P}\neq 0\}.$
For any function $z\in F$,
@@
is called the principal divisor of $z$. Given two divisors $G_{1}$ and $G_{2}$, we write $G_{1}\leq G_{2}$ if $n_{P}(G_{1})\leq n_{P}(G_{2})$ for all places $P$. The Riemann-Roch space associated with $G$ is defined as
@@
This space is a vector space over $\mathbb{F}_{q}$ with dimension given by the Riemann-Roch theorem:
@@
[43, Theorem 1.4.17] Let $\ell(G)$ denote the dimension of $\mathcal{L}(G)$ over $\mathbb{F}_{q}$. Then
@@ II-B Extensions of function fields
An algebraic function field $F^{\prime}/\mathbb{F}_{q^{m}}$ is called an algebraic extension of $F/\mathbb{F}_{q}$ if $F^{\prime}/F$ is an algebraic field extension with $F^{\prime}:F]<\infty$ and $m\geq 1$. A place $P^{\prime}$ of $F^{\prime}$ is said lying over a place $P$ of $F$ if $P\subseteq P^{\prime}$, denoted by $P^{\prime}|P$. The ramification index $e(P^{\prime}|P)$ is defined as the integer satisfying
@@
for all $x\in F$, and the relative degree of $P^{\prime}$ over $P$ is given by
@@
For any place $P$ of $F$, there exist only finitely many places $P^{\prime}$ of $F^{\prime}$ such that $P^{\prime}|P$. The conorm of $P$ is defined as
@@
Consequently, for any non-zero function $z\in F$, the principal divisor of $z$ in $F^{\prime}$ satisfies
@@
[43, Theorem 3.3.11] Let $P$ be a place of $F$, and let $P_{1},\ldots,P_{n}$ be all places of $F^{\prime}$ lying over $P$. Denote $e_{i}=e(P_{i}|P)$ and $f_{i}=f(P_{i}|P)$ for $i=1,\ldots,n$. Then
@@
If $F^{\prime}:F]=n$ and there exist exactly $n$ distinct places lying over $P$, then $P$ splits completely in $F^{\prime}/F$. In this case, $e_{i}=f_{i}=1$ for all $i$. If $e(P^{\prime}|P)>1$ for some $P^{\prime}$, then $P$ is ramified in $F^{\prime}/F$. Moreover, if there exists a place $P^{\prime}$ with $e(P^{\prime}|P)=n$, then $P$ is totally ramified in $F^{\prime}/F$.
II-C Galois Extensions
An extension $F^{\prime}/\mathbb{F}_{q^{m}}$ of a function field $F/\mathbb{F}_{q}$ is Galois if $F^{\prime}/F$ is a finite Galois extension. Specifically, the automorphism group
[…]
has order $F^{\prime}:F]$. This group is called the Galois group and write $\text{Gal}(F^{\prime}/F):=\text{Aut}(F^{\prime}/F)$.
For any place $P$ of $F/\mathbb{F}_{q}$, the Galois group $\text{Gal}(F^{\prime}/F)$ acts transitively on the set of places
[…]
via $\sigma(P^{\prime})=\{\sigma(x)\mid x\in P^{\prime}\}$. For any place $P^{\prime}$ of $F^{\prime}/\mathbb{F}_{q^{m}}$, the valuation corresponding to the place $\sigma(P^{\prime})$ is given by
$v_{\sigma(P^{\prime})}(y)=v_{P^{\prime}}(\sigma^{-1}(y)),\quad y\in F^{\prime}.$